Which one of the following suggests that the data set is approximately normal? O A. A data set with Q1 = 1330, Q3-2940, and s= 2440. O B. A data set with Q1 = 105, Q3-270, and s= 33. C. A data set with Q1 14, Q3-68, and s= 41 0 D. A data set with Q1-22, Q3-73, and s-21
Answer #1

Ans:

We know that for normal distribution:

IQR=Q3-Q1=1.349*s

Q1 Q3 s IQR=Q3-Q1 1.349*s
A 1330 2940 2440 1610 3291.56
B 105 270 33 165 44.517
C 14 68 41 54 55.309
D 2.2 7.3 2.1 5.1 2.8329

Option C is correct,as IQR is approximately 1.349*s